Scorecard Measures, Strategy Translation At the end of 20x1, Mejorar Company implemented a low-cost strategy to improve its competitive position. Its objective was to become the low-cost producer in its industry. A Balanced Scorecard was developed to guide the company toward this objective. To lower costs, Mejorar undertook a number of improvement activities such as JIT production, total quality management, and activity-based management. Now, after two years of operation, the president of Mejorar wants some assessment of the achievements. To help provide this assessment, the following information on one product has been gathered: 20x1 20x3 Theoretical annual capacity* 249,600 249,600 Actual production** 208,000 234,000 Market size (in units sold) 1,300,000 1,300,000 Production hours available (40 workers) 104,000 104,000 Very satisfied customers 62,400 140,400 Actual cost per unit $320 $256 Days of inventory 16 8 Number of defective units 12,480 7,020 Total worker suggestions 120 320 Hours of training 260 1,040 Selling price per unit $190 $190 Number of new customers 5,200 25,740 *Amount that could be produced given the available production hours; everything produced is sold. **Amount that was produced given the available production hours. Required: 1. Compute the following measures for 20x1 and 20x3: a. Actual velocity and cycle time. If required, round your answers to two decimal places. 20x1 20x3 Actual velocity per hour per hour Actual cycle time minutes minutes b. Percentage of total revenue from new customers (assume one unit per customer). Round your percentage answer to one decimal place (for example, enter 4.628% as "4.6"). 20x1 % 20x3 % c. Percentage of very satisfied customers (assume each customer purchases one unit). 20x1 % 20x3 % d. Market share. 20x1 % 20x3 % e. Percentage change in actual product cost (for 20x3 only). Enter a decrease as a negative value. 20x3 % f. Percentage change in days of inventory (for 20x3 only). Enter a decrease as a negative value. 20x3 % g. Defective units as a percentage of total units produced. Round your answers to two decimal places (for example, enter 4.628% as "4.63"). 20x1 % 20x3 % h. Total hours of training. 20x1 hours 20x3 hours i. Suggestions per production worker. Round your answers to one decimal place. 20x1 per production worker 20x3 per production worker j. Total revenue. 20x1 $ 20x3 $ k. Number of new customers. 20x1 20x3 2. Match the appropriate strategic objective to the measures given below. Strategic Objective Measure Unit cost reduction (%) Percentage of new revenues Revenue Percentage of very satisfied customers Market share Number of new customers Cycle time/Velocity Percentage defects Days of inventory Training hours Suggestions per worker
Answer #1

Solution:

1)

a)

Cycle time=Time between two successive deliveries

Cycle time=Total time/Actual production

Cycle time(2014)=104000/208000=0.5 hours=30 min

Cycle time(2016)=104000/234000=0.44 hours=26.67 min

Velocity=Number of units produced per hour

Velocity(2014)=208000/104000=2 units/hour

Velocity(2014)=234000/104000=2.25 units/hour

b)

Assuming 1 unit per consumer

Total customer=Total unit sold=Actual production

Percentage of revenue from new customers= New customers/Actual production

Percentage of revenue from new customers(2014)=5,200/208000 *100=2.5%

Percentage of revenue from new customers(2016)=25,740 /234000 * 100= 1%

c)

Assuming 1 unit per consumer

Total customer=Total unit sold=Actual production

Percentage of revenue from new customers= New customers/Actual production

Percentage of revenue from very satisfied customer(2014)=62,400/208000 *100= 30%

Percentage of revenue from new customers(2016)=140400/234000 * 100 =60%

d)

Market Share=Actual production/ Market Size

Market Share(2014)=208000/1300000 * 100=16%

Market Share(2016)=234000/1300000 * 100=18%

Note: Please re-post remaining questions. Happy to assist you !!